July 7, 2021 Calorimeter

From GlueXWiki
Jump to: navigation, search

Meeting Day/Time: every other WEDNESDAY at 11:00 a.m. JLab time

  1. To join via a Web Browser, go to the page [1] https://bluejeans.com/907185247.
  2. To join via Polycom room system go to the IP Address: (bjn.vc) and enter the meeting ID: 907185247.
  3. To join via phone, use one of the following numbers and the Conference ID: 907185247.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
  5. More information on connecting to bluejeans is available.

Participant Direct Lines

  • JLab Phone in CC L207 is 757-269-7084 (usual room)
  • JLab Phone: in CC F326 is 757-269-6460
  • Phone in the Regina Video-conference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947


  1. FCAL HDFCAL log book
  2. BCAL HDBCAL log book
  3. Calibrations: 2020 Data Production; RunPeriod-2019-11 Validation; Offline Monitoring Data Validation

Goals for Calorimetry Group

  1. Determine preliminary photon reconstruction efficiencies as a function of E, phi and theta in data and simulation with a point-to-point precision of at least 5%.
  2. Measure systematics of pi0/eta mass calibration as a function of detector position to a precision of at least 5 MeV.
  3. Demonstrate agreement of photon reconstruction efficiency and resolution between data and simulation as a function of E, phi and theta to within 5%.

Action Items

  1. Short term CALWG Items
  2. Long term items
  3. Calorimeter work packages

Tentative Agenda

  1. Announcements
  2. Run Preparations
    1. Adding failure modes to Hydra training
    2. FCAL: From Mark Dalton: "The main points are that the communications issues are probably all solved but that 5 bases will probably need to be replaced." See Log Entry 3883826
    3. BCAL
    4. CCAL
  3. Calibration
  4. Monitoring
    1. Detector Timing: From Mark Dalton: Monitoring launch version 33 shows good timing in all the timing detectors. See Log Entry 3883736
    2. BCAL LED
    3. BCAL (offline)
    4. FCAL (offline)
  5. Reconstruction
    1. Churamani: Brief Update from Spring18 data set regarding BCAL efficiency[2]
    2. Photon Reconstruction Task Force
  6. Simulations
  7. Action Items
  8. Any other business


Attendees: Elton, Tolga, Sean, Kishan, Varun, Zisis, Churamani, Jon, Thomas, Karthik, Sasha, Malte

  1. Announcements:
    • The run is on the horizon. See more below.
  2. Run Preparations
    1. Adding failure modes to Hydra training: Thomas and Kishan (EPSCI group) joined us. KIshan will be helping Thomas in improving the Hydra AI system, which we use to monitor the detectors. Thomas described the work done on the CDC: the AI can search for "bracket gate", single board failure events, etc. Thomas asked for CALWG expert help in improving "bad event" detection in the calorimeters: see how they appear in occupancy plots and classify them ("good", "bad", "hot channel", etc.). Elton mentioned that some BCAL channels get hot due to pedestal drift; these may need to be updated more frequently. Elton suggested that we also add 1D history to the Hydra monitoring, to complement the 2D ones which rescale.
    2. FCAL: From Mark Dalton: "The main points are that the communications issues are probably all solved but that 5 bases will probably need to be replaced." See Log Entry 3883826 Zisis summarized Mark's work on the FCAL; see log entry above for details. Procedure: i) Mark checked the EPICS db for the actual CAN IDs and block positions (pre replacement). ii) 209 bases were replaced. iii) Mark tracked the CAN IDs from the base serial number in JInventory. Results: i) he found duplicate entries and typos. ii) The past procedure did not remove CAN IDs from the base previous location. iii) Set up a dummy CAN ID of 7000000 when there is no CAN ID in JInventory. Summary: 5 bases need to be replaced (3 have vmon=0, 1 has no communication and 1 has vmon=1211 (set value is 1658).
    3. BCAL: Plans need to be made to take LED data before the run starts. Varun will coordinate with Mark.
    4. CCAL: Sasha reported on the activities: new dividers being installed, new LEMO connectors; broken strips being replaced; (uniform) torque wrench being used on the PMT plate; optical grease applied (10-15% of the area looks cloudy and will be re-examined); the LMS system is being modified to accommodate the larger (1mm diameter) fibres. Sasha expects all hardware to be ready by the end of July in the TDEF building and the detector will then be moved to Hall D.
  3. Calibration: no report.
  4. Monitoring
    1. Detector Timing: see Mark's note above. Elton asked if Tegan is looking at MC offsets in thing. Sean replied that Tegan was not doing it yet [At the following HDGeant Meeting this was discussed and Tegan will extract Δt (BCAL-RF) info.] Elton reminded the group that in the BCAL the charged particles have different offsets from the neutral ones, and this differs also between data and MC. Sean mentioned that we can using existing MC to do these studies. [Later on, Mark provided spectra comparing apples to apples between data and MC. MC is off by 70ps. Charged data is off by 150 ps. The MC seems to have slightly better resolution for photons and slightly worse resolution for charged.]
    2. BCAL LED: no report. Plans to take LED data (see above)
    3. BCAL (offline): nothing to be done here as all monitoring batches have been completed.
    4. FCAL (offline): Sean and Zisis will check with Susan to implement the dead channels in ccdb for the MC.
  5. Reconstruction
    1. Churamani gave a brief update from Spring18 data set regarding BCAL efficiency, using the Tag-and-Probe (TnP) method on the η. He examined the U1 and S1 flags; the former appears less biased. Elston asked about the selection cuts and Jon remarked that the π+π-γ(γ) in the FCAL were for the ω. We can get a ρ and a junk γ (split off). Elton suggested to make 2D plots Mγ(γ)π+π- vs Mπ+π-. Jon showed an efficiency plot from his past work, where a dip is send at 16 deg (CDC and FDC support structures) which now Churamin's plots are showing it: a good sign.
    2. Elton requested that Photon Reconstruction Task Force business stays as a permanent item in our CALWG agenda. Karthik will also pick up the b1 reconstruction soon. Jon commented that the η is only good up to 25-30 deg in the BCAL before dying out, while the b1 is superior in this regard but may suffer from 2γ (π0 mass) biasing.
  6. Simulations: no report.
  7. Action Items: as it was getting late, we will review these at our next meeting.
  8. Any other business: none.